# Metric Distortion in Peer Selection

### v1

@@ 3.2.1 Impossibility Results
Analogously, if $|S\cap(A_{1}\cup A_{2})|\geq q$, we define the distance metric $d_{3}$ on $A$ by the following positions $x\in(-\infty,\infty)^{n}$: $x_{a}=0$ for every $a\in A_{1}$, $x_{a}=1$ for every $a\in A_{2}$, and $x_{a}=2$ for every $a\in A_{3}\cup A_{4}$. It is not hard to check that $d_{3}\ \rhd\succ$; see Figure 3.(b) for an illustration. Since $|S\cap(A_{2}\cup A_{3}\cup A_{4})|\leq(k-q)+|A_{2}|=q-1<q$, we have that $\text{SC}(S,a;d_{3})=2$ for every $a\in A_{3}\cup A_{4}$. Furthermore, since both $|A_{1}|<q$ and $|A_{2}|<q$, we have that $\text{SC}(S,a;d_{3})=1$ for every $a\in A_{1}\cup A_{2}$. If we consider an alternative committee $S^{\prime}\subseteq A_{3}\cup A_{4}\in{A\choose k}$, which exists due to $|A_{3}\cup A_{4}|=3q-k-2\geq q$, we have $\text{SC}(S^{\prime},a;d_{3})=0$ for every $a\in A_{3}\cup A_{4}$, $\text{SC}(S^{\prime},a;d_{3})=1$ for every $a\in A_{2}$, and $\text{SC}(S^{\prime},a;d_{3})=2$ for every $a\in A_{1}$. Thus, we obtain
@@
The lower bound provided in this theorem increases in $q$ and varies between $\frac{3}{2}+\frac{3}{2(k+1)}$ for $q=\frac{k}{2}+1$ and $2$ for $q=k$; Figure 3.(a) illustrates it for $k=100$ and $q$ between $51$ and $100$.
@@ 3.2.2 A Voting Rule for $k=2$
On an intuitive level, the goal is to select agents that are both close to each other and close to the median agent(s). In particular, it is not hard to see that the optimal committee always consists of two consecutive agents: For any committee of non-consecutive agents, replacing the most extreme agent among the selected one with another closer to the median cannot decrease the social cost.
@@
Let $\mathcal{E}=(A,k,\succ)$ with $A=[n]$ and $d$ be as in the statement and $S^{\prime}=\{s_{1},s_{2}\}\in{A\choose k}$ an arbitrary committee. We assume that $s_{1}<s_{2}$ w.l.o.g.. Let $i\in\{1,\ldots,\big{\lfloor}\frac{n}{2}\rfloor\}$ be a fixed agent. If $i\leq s_{1}<s_{2}\leq n-i+1$, we have that the cost of the committee for agents $i$ and $n-i+1$ is at least
@@
In the remainder of this section, we establish tight distortion bounds of $\frac{4}{3}$ and $2$ for odd and even values of $n$, respectively. To derive these bounds, we frequently utilize the inequalities above to bound the social cost of the optimal committee.
@@
For the lower bound, on the other hand, we consider $n=5$ agents and two metrics illustrated in Figure 5. If a rule selects $\{1,2\}$, under metric $d_{1}$ the agents’ costs, from left to right, are $1$, $1$, $2$, $4$, and $4$, accounting for a total cost of $12$, while the optimal committee $\{4,5\}$ induces a social cost of $4+3+2+0+0=9$. Similarly, it is not hard to check that if a rule selects $\{2,3\}$, $\{3,4\}$, or $\{4,5\}$, then the social costs under metric $d_{2}$ are $8$, $8$, and $10$, respectively, while the optimal committee $\{1,2\}$ induces a social cost of $6$. In all cases, the ratio is at least $\frac{4}{3}$, so the bound follows.
@@
For each $i\leq\frac{n-1}{2}$, the joint cost of $S$ for agents $i$ and $n-i+1$ is given by
@@
11 11 $\succ_{1}2\succ_{1}3\succ_{1}4\succ_{1}5,$
22 22 $\succ_{2}1\succ_{2}3\succ_{2}4\succ_{2}5,$
33 33 $\succ_{3}2\succ_{3}1\succ_{3}4\succ_{3}5,$
44 44 $\succ_{4}5\succ_{4}3\succ_{4}2\succ_{4}1,$
55 55 $\succ_{5}4\succ_{5}3\succ_{5}2\succ_{5}1,$
@@
Suppose first that $S=\{1,2\}$. We take the distance metric $d_{1}$ on $A$ given by positions $x_{1}=0$, $x_{2}=1$, $x_{3}=2$, and $x_{4}=x_{5}=4$. It is not hard to check that $d_{1}\ \rhd\succ$; see Figure 5 for an illustration. Since $\text{SC}(\{1,2\},A;d_{1})=12$, and $\text{SC}(\{4,5\},A;d_{1})=9$, we obtain
@@
If $S\in\{\{2,3\},\{3,4\},\{4,5\}\}$, we consider the distance metric $d_{2}$ on $A$ given by positions $x_{1}=x_{2}=0$, $x_{3}=1$, $x_{4}=2$, and $x_{5}=3$. It is not hard to check that $d_{2}\ \rhd\succ$; see Figure 5 for an illustration. Since $\text{SC}(\{2,3\},A;d_{2})=\text{SC}(\{3,4\},A;d_{2})=8$ and $\text{SC}(\{4,5\},A;d_{2})=10$, whereas $\text{SC}(\{1,2\},A;d_{2})=6$, we obtain
@@
For an even number of agents $n$, the voting rule that selects the two median agents achieves a distortion of $2$ for utilitarian $2$-cost. Moreover, there exists $n\in\mathbb{N}$ such that, for every $(n,2)$-voting rule $f$, we have $\mathrm{dist}(f)\geq 2$ for utilitarian $2$-cost.
@@
11 11 $\succ_{1}2\succ_{1}3\succ_{1}4,$
22 22 $\succ_{2}1\succ_{2}3\succ_{2}4,$
33 33 $\succ_{3}4\succ_{3}2\succ_{3}1,$
44 44 $\succ_{4}3\succ_{4}2\succ_{4}1,$
@@
Suppose first that $S=\{3,4\}$. We take the distance metric $d_{1}$ on $A$ given by positions $x_{1}=x_{2}=0$, $x_{3}=1$, and $x_{4}=2$. It is not hard to check that $d_{1}\ \rhd\succ$; see Figure 6 for an illustration. Since $\text{SC}(\{3,4\},A;d_{1})=6$, and $\text{SC}(\{1,2\},A;d_{1})=3$, we obtain
@@
If $S=\{1,2\}$, we consider the distance metric $d_{2}$ on $A$ given by positions $x_{1}=0$, $x_{2}=1$, and $x_{3}=x_{4}=2$. It is not hard to check that $d_{2}\ \rhd\succ$; see Figure 6 for an illustration. Since $\text{SC}(\{1,2\},A;d_{2})=6$ and $\text{SC}(\{3,4\},A;d_{2})=3$, we obtain
@@
Finally, if $S=\{2,3\}$, we consider the distance metric $d_{3}$ on $A$ given by positions $x_{1}=x_{2}=0$ and $x_{3}=x_{4}=2$. It is not hard to check that $d_{3}\ \rhd\succ$; see Figure 6 for an illustration. Since $\text{SC}(\{2,3\},A;d_{3})=8$ and $\text{SC}(\{1,2\},A;d_{3})=\text{SC}(\{3,4\},A;d_{3})=4$, we obtain
@@
Metric $d_{1}$1,2121,21 , 23333444411111111Metric $d_{2}$111122223,4343,43 , 411111111Metric $d_{3}$1,2121,21 , 23,4343,43 , 42222
@@ 4.1 Warm-up: Selecting a Single Candidate
For every $n\in\mathbb{N}$, any $(n,1)$-voting rule has distortion $2$ for egalitarian social cost. There exists $n\in\mathbb{N}$ such that, for every $(n,1)$-voting rule $f$, $\mathrm{dist}(f)\geq 2$ for egalitarian social cost.
@@
Combining eqs. 14 and 15, we directly obtain that $\mathrm{dist}(f)\leq 2$.
@@
11 11 $\succ_{1}2\succ_{1}3\succ_{1}4,$
22 22 $\succ_{2}1\succ_{2}3\succ_{2}4,$
33 33 $\succ_{3}4\succ_{3}2\succ_{3}1,$
44 44 $\succ_{4}3\succ_{4}2\succ_{4}1,$
@@ 4.2 Egalitarian Additive Social Cost
We now distinguish two cases. If $a$ has at least as many agents in $S^{\prime}$ weakly to its left as strictly to its right; i.e., |$|\{s\in S^{\prime}\mid s\leq a\}|\geq|\{s\in S^{\prime}\mid s>a\}|$, then
@@
a contradiction to eq. 16. Analogously, if |$|\{s\in S^{\prime}\mid s\leq a\}|<|\{s\in S^{\prime}\mid s>a\}|$, then
@@
where we used the assumption $d\big{(}1,\big{\lfloor}\frac{k}{2}\big{\rfloor}\big{)}\leq d/2$ and the fact that $d(1,1)=0$ for the inequality. From Lemma 6 and eq. 17 we know that $\text{SC}(S^{\prime},A;d)\geq\frac{kd(1,n)}{2}$ for any $S^{\prime}\in{A\choose k}$, so we obtain
@@
where the inequality follows from the assumption $\sum_{s=2}^{\lfloor k/2\rfloor}d(1,s)\leq\frac{k-2-k\bmod 2}{k-1}\cdot\frac{kd(1,n)}{4}$ and the fact that $d(1,1)=0$. From Lemma 6 and eq. 17 we know that $\text{SC}(S^{\prime},A;d)\geq\frac{kd(1,n)}{2}$ for any $S^{\prime}\in{A\choose k}$, so we obtain
@@
Let now $S^{\prime}\in{A\setminus\{1\}\choose k-1}$ be any set of $k-1$ agents without $1$. Since $\big{\{}2,\ldots,\big{\lfloor}\frac{k}{2}\big{\rfloor}\big{\}}$ are the closest agents to $1$, we know that $\frac{1}{k-1}\sum_{s\in S^{\prime}}d(1,s)\geq\frac{1}{\lfloor k/2\rfloor-1}\sum_{s=2}^{\lfloor k/2\rfloor}d(1,s)$. Rearranging this expression and using our assumption once again, we obtain
@@
where we used eq. 18 for the last inequality. For any committee $S^{\prime}\in{A\choose k}$, this implies that $\text{SC}(S^{\prime},1;d)\geq\frac{kd(1,n)}{2}$, eq. 17 implies that $\text{SC}(S^{\prime},1;d)\geq\text{SC}(S^{\prime},n;d)$, and Lemma 6 implies that $\text{SC}(S^{\prime},A;d)=\text{SC}(S^{\prime},1;d)$. Therefore,
@@
Indeed, the first inequality follows from eq. 19 and the fact that $\text{SC}(S^{\prime},A;d)\geq\frac{kd(1,n)}{2}$ for every $S^{\prime}\in{A\choose k}$ due to eq. 17, the third equality from the definition of $f$, the second inequality from simple bounds on $d(1,s)$ for different values of $s$, and the last inequality from eq. 18. The other equalities come from simple calculations. Since the last expression again corresponds to the expression in the statement, we conclude.
@@ 4.3 Egalitarian $q$-Cost
    $b\succ_{a}c$ whenever $a\in A_{i},b\in A_{1},c\in A_{j}$ with $|i-1|=|i-j|$ for some $i,j\in[p]$;
@@
    $b\succ_{a}c$ whenever $a\in A_{i},b\in A_{p+1},c\in A_{j}$ with $|i-(p+1)|=|i-j|$ for some $i,j\in\{2,\ldots,p+1\}$;
@@
Let $n,k,q\in\mathbb{N}$ be such that $n\geq k\geq 2$ and $q>\frac{k}{3}$. If $q>\frac{k}{2}$, any $(n,k)$-voting rule has distortion $2$ for egalitarian $q$-cost. If $q>\frac{k}{3}$, $k$-Extremes has distortion $2$ for egalitarian $q$-cost. For every $k,q\in\mathbb{N}$ with $q>\frac{k}{3}\geq 1$, there exists $n\in\mathbb{N}$ with $n\geq k$ such that, for every $(n,k)$-voting rule $f$, $\mathrm{dist}(f)\geq 2$.
@@
Combining eqs. 20 and 21, we directly obtain that $\mathrm{dist}(f)\leq 2$.
@@
We first tackle two simple cases. If $a^{*}<q$, i.e., there are less than $q$ agents between $1$ and $a^{*}$, then for any committee $S^{\prime}\in{A\choose k}$ we have $\text{SC}(S^{\prime},A;d)\geq\text{SC}(S^{\prime},1;d)\geq d(1,a^{*}+1)\geq\frac{d(1,n)}{2}$, where the second inequality follows from eq. 22. Since $\text{SC}(S^{\prime},A;d)\leq d(1,n)$ holds for any committee $S^{\prime}\in{A\choose k}$, we know that in particular $\text{SC}(S,A;d)\leq d(1,n)$ and thus $\mathrm{dist}(f)\leq 2$. Similarly, if $n-a^{*}<q$, i.e., there are less than $q$ agents between $a^{*}+1$ and $n$, then for any committee $S^{\prime}\in{A\choose k}$ we have $\text{SC}(S^{\prime},A;d)\geq\text{SC}(S^{\prime},1;d)\geq d(a^{*},n)\geq\frac{d(1,n)}{2}$, where the second inequality follows from eq. 22. As before, $\mathrm{dist}(f)\leq 2$ thus follows directly.
@@
We claim that, for every $S^{\prime}\in{A\choose k}$, we have $\text{SC}(S^{\prime},A;d)\geq\frac{d^{*}}{2}$. Together with eq. 23, this would immediately imply $\mathrm{dist}(f)\leq 2$ and conclude the proof. To prove this fact, suppose for the sake of contradiction that $\text{SC}(S^{\prime},A;d)<\frac{d^{*}}{2}$ for some $S^{\prime}\in{A\choose k}$. This is equivalent to the fact that

### v2

@@ 3.2.1 Impossibility Results
Analogously, if $|S\cap(A_{1}\cup A_{2})|\geq q$, we define the distance metric $d_{3}$ on $A$ by the following positions $x\in(-\infty,\infty)^{n}$: $x_{a}=0$ for every $a\in A_{1}$, $x_{a}=1$ for every $a\in A_{2}$, and $x_{a}=2$ for every $a\in A_{3}\cup A_{4}$. Note that $d_{3}\ \rhd\succ$; see Figure 4.(b) for an illustration. Since $|S\cap(A_{2}\cup A_{3}\cup A_{4})|\leq(k-q)+|A_{2}|=q-1<q$, we have that $\text{SC}(S,a;d_{3})=2$ for every $a\in A_{3}\cup A_{4}$. Furthermore, since both $|A_{1}|<q$ and $|A_{2}|<q$, we have that $\text{SC}(S,a;d_{3})=1$ for every $a\in A_{1}\cup A_{2}$. If we consider an alternative committee $S^{\prime}\subseteq A_{3}\cup A_{4}\in{A\choose k}$, which exists due to $|A_{3}\cup A_{4}|=3q-k-2\geq q$, we have $\text{SC}(S^{\prime},a;d_{3})=0$ for every $a\in A_{3}\cup A_{4}$, $\text{SC}(S^{\prime},a;d_{3})=1$ for every $a\in A_{2}$, and $\text{SC}(S^{\prime},a;d_{3})=2$ for every $a\in A_{1}$. Thus, we obtain
@@
The lower bound provided in this theorem increases in $q$ and varies between $\frac{3}{2}+\frac{3}{2(k+1)}$ for $q=\frac{k}{2}+1$ and 22 for $q=k$; Figure 4.(a) illustrates it for $k=100$ and $q$ between $51$ and $100$.
@@ 3.2.2 A Voting Rule for $k=2$
On an intuitive level, the goal is to select agents that are both close to each other and close to the median agent(s). In particular, note that the optimal committee always consists of two consecutive agents: For any committee of non-consecutive agents, replacing the most extreme agent among the selected one with another closer to the median cannot decrease the social cost.
@@
Let $\mathcal{E}=(A,k,\succ)$ with $A=[n]$ and $d$ be as in the statement and $S^{\prime}=\{s_{1},s_{2}\}\in{A\choose k}$ an arbitrary committee. We assume that $s_{1}<s_{2}$ w.l.o.g.. Let $i\in\{1,\ldots,\big\lfloor\frac{n}{2}\rfloor\}$ be a fixed agent. If $i\leq s_{1}<s_{2}\leq n-i+1$, we have that the cost of the committee for agents $i$ and ni+1n-i+1 is at least
@@
In the remainder of this section, we establish tight distortion bounds of $\frac{4}{3}$ and 22 for odd and even values of $n$, respectively. To derive these bounds, we frequently utilize the inequalities above to bound the social cost of the optimal committee.
@@
For the lower bound, on the other hand, we consider $n=5$ agents and two metrics illustrated in Figure 6. If a rule selects $\{1,2\}$, under metric $d_{1}$ the agents’ costs, from left to right, are $1$, 11, 22, 44, and $4$, accounting for a total cost of $12$, while the optimal committee $\{4,5\}$ induces a social cost of $4+3+2+0+0=9$. Similarly, note that if a rule selects $\{2,3\}$, $\{3,4\}$, or $\{4,5\}$, then the social costs under metric $d_{2}$ are $8$, 88, and $10$, respectively, while the optimal committee $\{1,2\}$ induces a social cost of $6$. In all cases, the ratio is at least $\frac{4}{3}$, so the bound follows.
@@
For each $i\leq\frac{n-1}{2}$, the joint cost of $S$ for agents $i$ and ni+1n-i+1 is given by
@@
$1$ $\succ_{1}2\succ_{1}3\succ_{1}4\succ_{1}5,$
$2$ $\succ_{2}1\succ_{2}3\succ_{2}4\succ_{2}5,$
$3$ $\succ_{3}2\succ_{3}1\succ_{3}4\succ_{3}5,$
$4$ $\succ_{4}5\succ_{4}3\succ_{4}2\succ_{4}1,$
$5$ $\succ_{5}4\succ_{5}3\succ_{5}2\succ_{5}1,$
@@
Suppose first that $S=\{1,2\}$. We take the distance metric $d_{1}$ on $A$ given by positions $x_{1}=0$, $x_{2}=1$, $x_{3}=2$, and $x_{4}=x_{5}=4$. Note that $d_{1}\ \rhd\succ$; see Figure 6 for an illustration. Since $\text{SC}(\{1,2\},A;d_{1})=12$, and $\text{SC}(\{4,5\},A;d_{1})=9$, we obtain
@@
If $S\in\{\{2,3\},\{3,4\},\{4,5\}\}$, we consider the distance metric $d_{2}$ on $A$ given by positions $x_{1}=x_{2}=0$, $x_{3}=1$, $x_{4}=2$, and $x_{5}=3$. Note that $d_{2}\ \rhd\succ$; see Figure 6 for an illustration. Since $\text{SC}(\{2,3\},A;d_{2})=\text{SC}(\{3,4\},A;d_{2})=8$ and $\text{SC}(\{4,5\},A;d_{2})=10$, whereas $\text{SC}(\{1,2\},A;d_{2})=6$, we obtain
@@
For an even number of agents $n$, the voting rule that selects the two median agents achieves a distortion of 22 for utilitarian $2$-cost. Moreover, there exists $n\in\mathbb{N}$ such that, for every $(n,2)$-voting rule $f$, we have $\mathrm{dist}(f)\geq 2$ for utilitarian $2$-cost.
@@
$1$ $\succ_{1}2\succ_{1}3\succ_{1}4,$
$2$ $\succ_{2}1\succ_{2}3\succ_{2}4,$
$3$ $\succ_{3}4\succ_{3}2\succ_{3}1,$
$4$ $\succ_{4}3\succ_{4}2\succ_{4}1,$
@@
Suppose first that $S=\{3,4\}$. We take the distance metric $d_{1}$ on $A$ given by positions $x_{1}=x_{2}=0$, $x_{3}=1$, and $x_{4}=2$. Note that $d_{1}\ \rhd\succ$; see Figure 7 for an illustration. Since $\text{SC}(\{3,4\},A;d_{1})=6$, and $\text{SC}(\{1,2\},A;d_{1})=3$, we obtain
@@
If $S=\{1,2\}$, we consider the distance metric $d_{2}$ on $A$ given by positions $x_{1}=0$, $x_{2}=1$, and $x_{3}=x_{4}=2$. Note that $d_{2}\ \rhd\succ$; see Figure 7 for an illustration. Since $\text{SC}(\{1,2\},A;d_{2})=6$ and $\text{SC}(\{3,4\},A;d_{2})=3$, we obtain
@@
Finally, if $S=\{2,3\}$, we consider the distance metric $d_{3}$ on $A$ given by positions $x_{1}=x_{2}=0$ and $x_{3}=x_{4}=2$. Note that $d_{3}\ \rhd\succ$; see Figure 7 for an illustration. Since $\text{SC}(\{2,3\},A;d_{3})=8$ and $\text{SC}(\{1,2\},A;d_{3})=\text{SC}(\{3,4\},A;d_{3})=4$, we obtain
@@
Metric $d_{1}$1,21,233441111Metric $d_{2}$11223,43,41111Metric $d_{3}$1,21,23,43,422
@@ 4.1 Warm-up: Selecting a Single Candidate
For every $n\in\mathbb{N}$, any $(n,1)$-voting rule has distortion 22 for egalitarian social cost. There exists $n\in\mathbb{N}$ such that, for every $(n,1)$-voting rule $f$, $\mathrm{dist}(f)\geq 2$ for egalitarian social cost.
@@
Combining inequalities 18 and 19, we directly obtain that $\mathrm{dist}(f)\leq 2$.
@@
$1$ $\succ_{1}2\succ_{1}3\succ_{1}4,$
$2$ $\succ_{2}1\succ_{2}3\succ_{2}4,$
$3$ $\succ_{3}4\succ_{3}2\succ_{3}1,$
$4$ $\succ_{4}3\succ_{4}2\succ_{4}1,$
@@ 4.2 Egalitarian Additive Social Cost
We now distinguish two cases. If $a$ has at least as many agents in $S^{\prime}$ weakly to its left as strictly to its right; i.e., $|\{s\in S^{\prime}\mid s\leq a\}|\geq|\{s\in S^{\prime}\mid s>a\}|$, then
@@
a contradiction to inequality 20. Analogously, if $|\{s\in S^{\prime}\mid s\leq a\}|<|\{s\in S^{\prime}\mid s>a\}|$, then
@@
where we used the assumption $d\big(1,\big\lfloor\frac{k}{2}\big\rfloor\big)\leq d/2$ and the fact that $d(1,1)=0$ for the inequality. From Lemma 6 and inequality 21 we know that $\text{SC}(S^{\prime},A;d)\geq\frac{kd(1,n)}{2}$ for any $S^{\prime}\in{A\choose k}$, so we obtain
@@
where the inequality follows from the assumption $\sum_{s=2}^{\lfloor k/2\rfloor}d(1,s)\leq\frac{k-2-k\bmod 2}{k-1}\cdot\frac{kd(1,n)}{4}$ and the fact that $d(1,1)=0$. From Lemma 6 and inequality 21 we know that $\text{SC}(S^{\prime},A;d)\geq\frac{kd(1,n)}{2}$ for any $S^{\prime}\in{A\choose k}$, so we obtain
@@
Let now $S^{\prime}\in{A\setminus\{1\}\choose k-1}$ be any set of k1k-1 agents without $1$. Since $\big\{2,\ldots,\big\lfloor\frac{k}{2}\big\rfloor\big\}$ are the closest agents to $1$, we know that $\frac{1}{k-1}\sum_{s\in S^{\prime}}d(1,s)\geq\frac{1}{\lfloor k/2\rfloor-1}\sum_{s=2}^{\lfloor k/2\rfloor}d(1,s)$. Rearranging this expression and using our assumption once again, we obtain
@@
where we used inequality 22 for the last inequality. For any committee $S^{\prime}\in{A\choose k}$, this implies that $\text{SC}(S^{\prime},1;d)\geq\frac{kd(1,n)}{2}$, inequality 21 implies that $\text{SC}(S^{\prime},1;d)\geq\text{SC}(S^{\prime},n;d)$, and Lemma 6 implies that $\text{SC}(S^{\prime},A;d)=\text{SC}(S^{\prime},1;d)$. Therefore,
@@
Indeed, the first inequality follows from inequality 23 and the fact that $\text{SC}(S^{\prime},A;d)\geq\frac{kd(1,n)}{2}$ for every $S^{\prime}\in{A\choose k}$ due to inequality 21, the third equality from the definition of $f$, the second inequality from simple bounds on $d(1,s)$ for different values of $s$, and the last inequality from inequality 22. The other equalities come from simple calculations. Since the last expression again corresponds to the expression in the statement, we conclude.
@@ 4.3 Egalitarian $q$-Cost
    $b\succ_{a}c$ whenever $a\in A_{i},b\in A_{1},c\in A_{j}$ with |i1|=|ij||i-1|=|i-j| for some $i,j\in[p]$;
@@
    $b\succ_{a}c$ whenever $a\in A_{i},b\in A_{p+1},c\in A_{j}$ with |i(p+1)|=|ij||i-(p+1)|=|i-j| for some $i,j\in\{2,\ldots,p+1\}$;
@@
Let $n,k,q\in\mathbb{N}$ be such that $n\geq k\geq 2$ and $q>\frac{k}{3}$. If $q>\frac{k}{2}$, any $(n,k)$-voting rule has distortion 22 for egalitarian $q$-cost. If $q>\frac{k}{3}$, $k$-Extremes has distortion 22 for egalitarian $q$-cost. For every $k,q\in\mathbb{N}$ with $q>\frac{k}{3}\geq 1$, there exists $n\in\mathbb{N}$ with $n\geq k$ such that, for every $(n,k)$-voting rule $f$, $\mathrm{dist}(f)\geq 2$.
@@
Combining inequalities 24 and 25, we directly obtain that $\mathrm{dist}(f)\leq 2$.
@@
We first tackle two simple cases. If $a^{*}<q$, i.e., there are less than $q$ agents between $1$ and $a^{*}$, then for any committee $S^{\prime}\in{A\choose k}$ we have $\text{SC}(S^{\prime},A;d)\geq\text{SC}(S^{\prime},1;d)\geq d(1,a^{*}+1)\geq\frac{d(1,n)}{2}$, where the second inequality follows from inequality 26. Since $\text{SC}(S^{\prime},A;d)\leq d(1,n)$ holds for any committee $S^{\prime}\in{A\choose k}$, we know that in particular $\text{SC}(S,A;d)\leq d(1,n)$ and thus $\mathrm{dist}(f)\leq 2$. Similarly, if $n-a^{*}<q$, i.e., there are less than $q$ agents between $a^{*}+1$ and $n$, then for any committee $S^{\prime}\in{A\choose k}$ we have $\text{SC}(S^{\prime},A;d)\geq\text{SC}(S^{\prime},1;d)\geq d(a^{*},n)\geq\frac{d(1,n)}{2}$, where the second inequality follows from inequality 26. As before, $\mathrm{dist}(f)\leq 2$ thus follows directly.
@@
We claim that, for every $S^{\prime}\in{A\choose k}$, we have $\text{SC}(S^{\prime},A;d)\geq\frac{d^{*}}{2}$. Together with inequality 27, this would immediately imply $\mathrm{dist}(f)\leq 2$ and conclude the proof. To prove this fact, suppose for the sake of contradiction that $\text{SC}(S^{\prime},A;d)<\frac{d^{*}}{2}$ for some $S^{\prime}\in{A\choose k}$. This is equivalent to the fact that